# Quantum resolution limit of long-baseline imaging using distributed entanglement

### v1

@@ III.3 The state over multiple time bins
where $\rho_{1,AB,m}$ is a tensor product of $M$ states of which $M-1$ are vacuum and the remaining one is a one-photon state $\rho_{1,AB}$ in the $m$-th position:
$\rho_{1,AB,m}$ $=\rho_{0,AB}^{\otimes(m-1)}\otimes\rho_{1,AB}\otimes\rho_{0,AB}^{\otimes(M-m)}$ (29a)
$=\rho_{0,AB}^{\otimes(m-1)}\otimes\frac{1}{2}\sum_{s\in\{1,2\}}\left|\psi^{(s)}_{AB,K}\right\rangle\left\langle\psi^{(s)}_{AB,K}\right|\otimes\rho_{0,AB}^{\otimes(M-m)}$ (29b)
@@
The ket in Eq. (29c) is for a single photon from star $s$ arriving at time bin $m$ defined as
@@
$\left|\widetilde{0}_{\alpha}\right\rangle\equiv\bigotimes_{i=0}^{K-1}\bigotimes_{j=1}^{M}|0_{\alpha_{ji}}\rangle,\quad\alpha\in\{A,B\}$ (31)
@@
$\left|\widetilde{1}_{\alpha_{mq}}\right\rangle=a_{\alpha_{mq}}^{\dagger}\left|\widetilde{0}_{\alpha}\right\rangle,\quad\alpha\in\{A,B\}$ (32)
are the kets for vacuum across the $KM$ modes and one photon in spatio-temporal mode $mq$, respectively. The creation operator $a_{\alpha_{mq}}^{\dagger}$ produces a photon in the $mq$-th spatio-temporal mode at site $\alpha$. Since we are only considering zero photon states and single photon excitations, we can think of the state associated with each spatio-temporal mode at a given site as a single rail qubit, which is in $|0\rangle$ if there is no photon, and $|1\rangle$ if there is one photon.
In the next sections, we address the different stages towards interferometry of the incoming photons. First we go through the encoding process in which starlight information is mapped onto a quantum memory in the fashion of Khabiboulline et al. (2019a, b). In the decoding section, we show how the memory is mapped onto ancilla Bell pairs to transfer the information from one of the aperture sites to the other.
@@ IV Encoding photonic quantum state into quantum memories
The encoding step involves transferring the state of the photons to a quantum memory state described by sets of systems $\overline{A}$ and $\overline{B}$, labeled $\overline{\alpha}\in\{\overline{A},\overline{B}\}$, which start as a register, for a given spatial mode, of $\overline{M}=\log_{2}(M+1)$ qubits initialized in the state $|0\rangle$. We introduce indices $k\in\{1,\dots,\overline{M}\}$, $i\in\{0,\dots,K-1\}$ and $j\in\{1,\dots,M\}$ that account for the resources used throughout the protocol (memory qubits, CNOT gates and pre-shared Bell pairs, respectively), with the special case $i=q$ and $j=m$ corresponding to the spatio-temporal mode $mq$ where the photonic excitation is present. For both sites we write the state of the memory qubits as
@@
We then wish to load the photonic state on to the memory qubits. For this purpose, we employ a collection of CNOT gates from the photonic qubits to the memory ones such that a photonic excitation in the $i$th spatial mode in the $j$th time bin at a given site translates into excitations in the binary representation of $j$ in the memory qubits corresponding to spatial mode $i$ at the same location. These CNOT operations for site $\alpha$ can be denoted by $U_{\alpha_{ji},\overline{\alpha}_{ki}}$, where we identify the control and target qubits as the photonic qubit in the $i$-th spatial mode and $j$-th temporal mode (photonic qubit $\alpha_{ji}$) and the $k$-th qubit in the memory (memory qubit $\overline{\alpha}_{ki}$), respectively. The full operator acting on the state for all spatial and temporal modes at site $\alpha$ is
@@
The result from this mapping is thus an entangled state between the photon and the memory. The next step is to disentangle the two and separate the photons from this. This is done by measuring the photonic qubits (systems $A$ and $B$) in the $X$ basis |±$|\pm\rangle=(|0\rangle\pm|1\rangle)/\sqrt{2}$. Such measurement of the photonic modes will introduce a conditional phase in the state of the memory qubits. One way we can deal with this is by expanding the photonic portions of (40) into a sum of four $X$ basis tensor product states. First note that going from the $Z$ basis to the $X$ basis for two modes yields the following:
@@
$\quad\left.-|+_{A_{ji}}-_{B_{ji}}\rangle+|-_{A_{ji}}+_{B_{ji}}\rangle\right),$ (41)
@@
Thus from (44), we see that measuring the photonic qubits in the $X$ basis dis-entangles the photon-memory system, but gives a conditional sign flip in the memory state depending on whether we obtain odd (±,±)plus-or-minusplus-or-minus($\pm$,$\pm$)( ± , ± ) or even (±,)plus-or-minusminus-or-plus($\pm$,$\mp$)( ± , ∓ ) parity results for the qubits associated with the temporal and spatial modes $mq$ at the two sites. and the resulting memory state is
@@
Hhere the conditional sign is given by
@@
where $h_{A_{mq}}$ and $h_{B_{mq}}$ are the results of the $X$ basis measurements of the photonic qubits corresponding to the time bin $m$ and spatial mode $q$ at the two sites. We should mention that at this point, we do not yet know the temporal and spatial mode $mq$ of the incoming photon, and therefore cannot tell if $f(\textbf{h}_{mq})$ is $1$ or $-1$. Thus at this stage, we have to simply store the results of all the $X$ basis measurement results in a classical memory register for now. Once we go through the decoding stage and determine the temporal and spatial mode, then we will be able to look at these stored results and see if $f(\textbf{h}_{mq})$ is $1$ or $-1$, depending on whether the corresponding photonic qubit results have even or odd parity, respectively. Therefore, $h_{A_{mq}}$ and $h_{A_{mq}}$ are entries of a $2KM$-component array $h$ containing the results of all the $X$ basis measurement outcomes.
@@
The encoding of the arrival time is complete, nonetheless, we require the collapse of the superposition over index $q$ via a projective measurement, which is accomplished in the decoding step which we describe in the next section. The implementation of the encoding would require the interaction of the incoming photon with an atomic cavity, similar to the one described in Ref. Duan and Kimble (2004). As for the measurement of the photonic state in the $X$ basis, it has been suggested in Ref. Khabiboulline et al. (2019b) that it could be approximated by mixing the photon with ancillary states in an interferometer, and in principle, the larger the interferometer, the closest we could be to a deterministic $X$-basis measurement. A more careful study of this proposal and an alternative method for a deterministic $X$ basis measurement will be presented in a future publication.
@@ V Decoding
We now proceed with the decoding step where we reveal the time bins in which the photons arrive. The employed tool is an ancilla consisting of $K\overline{M}$ entangled pairs described by the sets of systems $C$ and $D$. This ancilla state has the form
$|\Phi_{CD}\rangle=\bigotimes_{i=0}^{K-1}\bigotimes_{k=1}^{\overline{M}}\left|\phi^{+}_{C_{ki}D_{ki}}\right\rangle.$ (50)
where $|\phi^{+}_{C_{ki}D_{ki}}\rangle=(|0_{C_{ki}},0_{D_{ki}}\rangle+|1_{C_{ki}},1_{D_{ki}}\rangle)/\sqrt{2}$. We will then employ a circuit that flip some or all of these into $|\phi^{-}_{C_{ki}D_{ki}}\rangle=(|0_{C_{ki}},0_{D_{ki}}\rangle-|1_{C_{ki}},1_{D_{ki}}\rangle)/\sqrt{2}$, if some of the memory qubits are in $|1\rangle$ due to an incoming photon in the original photonic system. The Bell states upon measurement in the $X$ basis, will yield ‘even’ or ‘odd’ parity results and reveal which of the Bell states underwent a flip and allow us to determine the spatio-temporal mode of the incoming photon, as we will shortly describe. This step is also known as ‘checking’ the parity.
@@
is the Bell state ancilla after the unitary $V$ has been applied. The Pauli operators $Z_{C_{ki}}$ and $Z_{D_{ki}}$ have the effect of flipping the sign of the relevant Bell pairs according to the rules (52a)-(52d) (by acting on qubit $ki$ of system $C$ and $D$), respectively. The power $\delta_{iq}$ means that this operator is acting only when $i=q$, and $w_{km}$ means it is happening only for the Bell pairs corresponding to the binary representation of the temporal mode number $m$. In short, if we have an incoming photon in temporal mode $m$, then in each term of our state, the Bell pairs corresponding to the spatio-temporal mode $mq$ undergo a flip, while the remaining Bell states are unflipped. On the other hand, the action of $V$ on the first term of (49) associated with the photonic vacuum of systems $A$ and $B$, will leave all the Bell pairs unaffected: $V|0_{\overline{A}},0_{\overline{B}}\rangle|\Phi_{CD}\rangle=|0_{\overline{A}},0_{\overline{B}}\rangle|\Phi_{CD}\rangle$.
Before performing the parity checks, we write the Bell states in terms of the $X$ basis:
@@
Having found the spatial mode, the remaining task is to determine if the photon is in the symmetric or anti-symmetric combination of mode $q$ collected at the two sites. Now, the memory qubits corresponding to the unflipped Bell states Now, the memory qubits corresponding to the unflipped Bell states (i.e. the ones found to have even parity) are all in $\left|0\right\rangle$. These contain no useful information and we can drop them going forward. This leaves only the memory qubits corresponding to the Bell pairs found to be in $|\phi^{-}\rangle$ (i.e. the ones for which we obtained odd parity results). Let $N_{m}\in\{1,...,\overline{M}\}$ be the number of such Bell pairs found to have flipped into $\left|\phi^{-}\right\rangle$ after the CZ gates step for a given time bin $m$ in which the photon arrived. We can now define subsystems $E$ and $F$ containing only the $2N_{m}$ remaining memory qubits, and we adopt the labeling $\gamma\in\{E,F\}$ for these new systems. Then we construct the reduced state consisting of $2N_{m}$ entangled qubits, and this has the form
@@
Note that we have defined the states (61) in terms of $x_{s}$. We will make the substitutions $x_{1}=\theta$ and $x_{2}=-\theta$ for our 2-point problem later.
@@
At this point, the task of measuring (61) in the (64) states is, in principle, a GHZ state measurement. This can be done by measuring each qubit in the $X$-basis and then considering the parities of the outcomes associated with the results for each pair of qubits at both sites corresponding to the same digit of the binary encoding of the spatio-temporal mode $mq$. We show in appendix B that an even number of odd parity results i.e. $|+_{E_{\mu q}},-_{F_{\mu q}}\rangle$ or $|-_{E_{\mu q}},+_{F_{\mu q}}\rangle$ means finding the state to be $|\zeta^{+}_{mq}\rangle$, and an odd number of odd parity results signifies obtaining $|\zeta^{-}_{EF,mq}\rangle$ as our measurement outcome. Putting together the pieces along with the two cases for $f(\textbf{h}_{mq})$, we obtain the conditional probabilities $p_{\pm}(x_{s})$ that given an incoming photon is in spatial mode $q$, it is in the symmetric or anti-symmetric combinations $|\phi_{AB,q}^{\pm}\rangle$ of the signals collected at the two sites
@@ VI Quantum limit of precision for estimating two-source separation
where $\psi(x)$ is the PSF, and $N$ is the total number of incoming photons during the integration time. Taking the combined PSF (13) for our system system of 2 hard apertures, and setting $\beta=\pi r/\sigma$, where $r=2\beta/\delta$ is the ratio between the mid-point separation $2\beta$ (between the two apertures) and the individual aperture size $\delta=2\pi/\sigma$, gives the following QFI
@@
Now, recall that $\sigma=2\pi/\delta$, with $\delta$ the size and $r=2\beta/\delta$. Also, the total number $N$ of photons collected during the integration time will be proportional to the area of the apertures $\delta^{2}$. So all in all, the $4N\pi^{2}/\sigma^{2}$ factor will scale as $\delta^{4}$. The $3r^{2}$ term in the QFI will scale as $\delta^{2}\beta^{2}$, whereas the other term will be proportional to $\delta^{4}$. This tells us that as $\delta$ gets larger (or equivalently as $r$ is small and closer to unity), we will get more information from the spatial mode resolution of the light within each aperture.

### v2

@@ III.3 Quantum state of the collected light for two telescopes over $M$ temporal modes
where $\rho_{1,AB,m}$ is a tensor product of $M$ states of which $M-1$ are vacuum and the remaining one is a one-photon state $\rho_{1,AB}$ in the $m$-th position (of the tensor product):
$\rho_{1,AB,m}$ $=$ $\rho_{0,AB}^{\otimes(m-1)}\otimes\rho_{1,AB}\otimes\rho_{0,AB}^{\otimes(M-m)}$ (30b)
$=$ $\rho_{0,AB}^{\otimes(m-1)}\otimes\frac{1}{2}\sum_{s\in\{1,2\}}\left|\psi^{(s)}_{AB,K}\right\rangle\left\langle\psi^{(s)}_{AB,K}\right|$
$\otimes\rho_{0,AB}^{\otimes(M-m)}$
@@
The ket in Eq. (30c) is that of a single photon from star $s$ arriving in temporal mode indexed $m$ (in the time interval $T$), and is defined as:
@@
$\left|\widetilde{0}_{\alpha}\right\rangle\equiv\bigotimes_{i=0}^{K-1}\bigotimes_{j=1}^{M}|0_{\alpha_{ji}}\rangle,\quad\alpha\in\{A,B\},$ (32)
@@
$\left|\widetilde{1}_{\alpha_{mq}}\right\rangle=a_{\alpha_{mq}}^{\dagger}\left|\widetilde{0}_{\alpha}\right\rangle,\quad\alpha\in\{A,B\},$ (33)
are the states corresponding to vacuum across the $KM$ modes and one photon in the spatio-temporal mode $(m,q)$, respectively. The creation operator $a_{\alpha_{mq}}^{\dagger}$ produces a photon in the $(m,q)$-th spatio-temporal mode at telescope site $\alpha$.
Since we are only considering zero and single photon states across the $M$ temporal modes, for the subsequent sections when we discuss photon-memory interactions for loading the science photons into quantum memories, we will often find it useful to think of the state associated with each spatio-temporal mode of the collected optical field at a given telescope site as a single rail photonic qubit, i.e., a qubit encoded in the span of $|0\rangle$ (vacuum state of that spatio-temporal mode) and $|1\rangle$ (the state of $1$ photon in that spatio-temporal mode).
In the subsequent Sections, we address the different stages towards entanglement-assisted (nonlocal) interferometry of the incoming photon field collected at the two telescopes across the $KM$ spatio-temporal modes described above. First we go through the encoding process in which the information in the science photons is mapped onto qubits of a quantum memory register in the fashion of Khabiboulline et al. (2019a, b). In the Section of decoding, we will show how the state of the memories are mapped onto ancilla Bell pairs for (teleported) transfer of the information from one of the aperture sites to the other, followed by a final measurement sequence that effectively realizes the desired projection (24) on the original (optical) state of all the $KM$ mode pairs at the two telescope sites.
@@ IV Encoding photonic quantum state into quantum memories
The encoding step involves transferring the state of the photons to a quantum memory state described by sets of systems $\overline{A}$ and $\overline{B}$, labeled $\overline{\alpha}\in\{\overline{A},\overline{B}\}$, which start as a register, for a given spatial mode, of $\overline{M}=\log_{2}(M+1)$ qubits initialized in the state $|0\rangle$. We introduce indices $k\in\{1,\dots,\overline{M}\}$, $i\in\{0,\dots,K-1\}$ and $j\in\{1,\dots,M\}$ that account for the resources used throughout the protocol (memory qubits, CNOT gates and pre-shared Bell pairs, respectively), with the special case $i=q$ and $j=m$ corresponding to the spatio-temporal mode $(m,q)$ where the photon is present. For both sites we write the state of the memory qubits as
@@
We then wish to load the photonic state on to the memory qubits. From hereon, when referring to one of the $KM$ optical modes, we will synonymously use photonic qubit to refer to the single-rail photonic qubit (of the respective mode). We first employ a collection of CNOT gates from the $KM$ photonic qubits to the $K{\bar{M}}$ memory qubits, such that a photonic excitation in the $i$th spatial mode of the $j$th temporal mode at a given telescope site ($\alpha$ or $\beta$) translates into excitation(s) in the binary representation of the integer $j$ into the memory qubits of the memory-register corresponding to spatial mode $i$ at that telescope location. These CNOT operations for site $\alpha$ can be denoted by $U_{\alpha_{ji},\overline{\alpha}_{ki}}$, where we identify the control and target qubits as the photonic qubit in the $i$-th spatial mode and $j$-th temporal mode (photonic qubit $\alpha_{ji}$) and the $k$-th qubit in the memory (memory qubit $\overline{\alpha}_{ki}$), respectively. The full unitary operator acting on the state for all spatial and temporal modes at site $\alpha$ is:
@@
The result from this mapping is thus an entangled state between the single-photon quantum state of the $2MK$ spatio-temporal modes of the collected field and the $2{\bar{M}}$ qubits of the two ${\bar{M}}$-qubit memory registers at the telescope sites. The next step is to disentangle the two and separate the photon from the memory. This is done by measuring the photonic qubits (systems $A$ and $B$) in the $X$ basis |±$|\pm\rangle=(|0\rangle\pm|1\rangle)/\sqrt{2}$. Such measurement of the photonic modes will introduce a conditional phase in the state of the memory qubits. One way we can deal with this is by expanding the photonic portions of (41) into a sum of four $X$ basis tensor product states. First note that going from the $Z$ basis to the $X$ basis for two modes yields the following:
@@
$\quad\left.|+_{A_{ji}}-_{B_{ji}}\rangle+|-_{A_{ji}}+_{B_{ji}}\rangle\right),$ (42)
@@
Thus from (45), we see that measuring the photonic qubits in the $X$ basis dis-entangles the photon-memory system, but gives a conditional sign flip in the memory state depending on whether we obtain odd (±,±)plus-or-minusplus-or-minus($\pm$,$\pm$)( ± , ± ) or even (±,)plus-or-minusminus-or-plus($\pm$,$\mp$)( ± , ∓ ) parity results for the qubits associated with the temporal and spatial modes $(m,q)$ at the two sites. and the resulting memory state is
@@
Here the conditional sign is given by
@@
where $h_{A_{mq}}$ and $h_{B_{mq}}$ are the results of the $X$ basis measurements of the photonic qubits corresponding to the time bin $m$ and spatial mode $q$ at the two sites.
We should mention that at this point, we do not yet know the temporal and spatial mode $(m,q)$ of the incoming photon, and therefore cannot tell if $f(\textbf{h}_{mq})$ is $1$ or $-1$. Thus at this stage, we have to simply store the results of all the $X$ basis measurement results in a classical memory register for now. Once we go through the decoding stage and determine the temporal and spatial mode, then we will be able to look at these stored results and see if $f(\textbf{h}_{mq})$ is $1$ or $-1$, depending on whether the corresponding photonic qubit results have even or odd parity, respectively. Therefore, $h_{A_{mq}}$ and $h_{B_{mq}}$ are entries of a $2KM$-component array $h$ containing the results of all the $X$ basis measurement outcomes.
@@
The encoding of the arrival time is complete, nonetheless, we require the collapse of the superposition over index $q$ via a projective measurement, which is accomplished in the decoding step which we describe in the next section. The implementation of the encoding would require the interaction of the incoming photon with an atomic cavity, similar to the one described in Ref. Duan and Kimble (2004). As for the measurement of a single-rail photonic qubit in the $X$ basis, it has been suggested in Ref. Khabiboulline et al. (2019b) that it could be approximated by mixing the photon with ancillary states in an interferometer, and in principle, the larger the interferometer, the closest we could be to a deterministic $X$-basis measurement. A more careful analysis of this proposal and an alternative method for an in-principle deterministic $X$ basis measurement on single-rail photonic qubits will be presented in future publications Sajjad2025b; Sajjad2025c.
@@ V Decoding
We now proceed with the decoding step where we reveal the time bins in which the photons arrive. The employed tool is an ancilla consisting of $K\overline{M}$ entangled pairs described by the sets of systems $C$ and $D$ This ancilla state has the form
$|\Phi_{CD}\rangle=\bigotimes_{i=0}^{K-1}\bigotimes_{k=1}^{\overline{M}}\left|\phi^{+}_{C_{ki}D_{ki}}\right\rangle,$ (51)
where $|\phi^{+}_{C_{ki}D_{ki}}\rangle=(|0_{C_{ki}},0_{D_{ki}}\rangle+|1_{C_{ki}},1_{D_{ki}}\rangle)/\sqrt{2}$. We will then employ a circuit that flips some or all of these into $|\phi^{-}_{C_{ki}D_{ki}}\rangle=(|0_{C_{ki}},0_{D_{ki}}\rangle-|1_{C_{ki}},1_{D_{ki}}\rangle)/\sqrt{2}$, if some of the memory qubits are in $|1\rangle$ due to an incoming photon in the original photonic system. The Bell states upon measurement in the $X$ basis, will yield ‘even’ or ‘odd’ parity results and reveal which of the Bell states underwent a flip and allow us to determine the spatio-temporal mode of the incoming photon, as we will shortly describe. This step is also known as ‘checking’ the parity.
@@
is the Bell state ancilla after the unitary $V$ has been applied. The Pauli operators $Z_{C_{ki}}$ and $Z_{D_{ki}}$ have the effect of flipping the sign of the relevant Bell pairs according to the rules (53a)-(53d) (by acting on qubit $(k,i)$ of system $C$ and $D$), respectively. The power $\delta_{iq}$ means that this operator is acting only when $i=q$, and $w_{km}$ means it is happening only for the Bell pairs corresponding to the binary representation of the temporal mode number $m$. In short, if we have an incoming photon in temporal mode $m$, then in each term of our state, the Bell pairs corresponding to the spatio-temporal mode $(m,q)$ undergo a flip, while the remaining Bell states are unflipped. On the other hand, the action of $V$ on the first term of (50) associated with the photonic vacuum of systems $A$ and $B$, will leave all the Bell pairs unaffected: $V|0_{\overline{A}},0_{\overline{B}}\rangle|\Phi_{CD}\rangle=|0_{\overline{A}},0_{\overline{B}}\rangle|\Phi_{CD}\rangle$.
We now need to determine which Bell pairs flipped. For this, note that expressing the flipped and unflipped Bell states in the $X$ basis, we obtain
@@
Having found the spatial mode, the remaining task is to determine if the photon is in the symmetric or anti-symmetric combination of mode $q$ collected at the two sites. Now, the memory qubits corresponding to the unflipped Bell states (i.e., the ones found to have even parity) are all in $\left|0\right\rangle$. These contain no useful information and we can drop them going forward. This leaves only the memory qubits corresponding to the Bell pairs found to be in $|\phi^{-}\rangle$ (i.e. the ones for which we obtained odd parity results). Let $N_{m}\in\{1,...,\overline{M}\}$ be the number of such Bell pairs found to have flipped into $\left|\phi^{-}\right\rangle$ after the CZ gates step for a given time bin $m$ in which the photon arrived. We can now define subsystems $E$ and $F$ containing only the $2N_{m}$ remaining memory qubits, and we adopt the labeling $\gamma\in\{E,F\}$ for these new systems. Then we construct the reduced state consisting of $2N_{m}$ entangled qubits, and this has the form
@@
Note that we have defined the states (62) in terms of $x_{s}$. We will make the substitutions $x_{1}=\theta$ and $x_{2}=-\theta$ for our 2-point-source separation-estimation problem later.
@@
At this point, the task of measuring (62) in the (65) states is, in principle, a GHZ state measurement. This can be done by measuring each qubit in the $X$-basis and then considering the parities of the outcomes associated with the results for each pair of qubits at both sites corresponding to the same digit of the binary encoding of the spatio-temporal mode $(m,q)$. We show in appendix B that an even number of odd parity results i.e. $|+_{E_{\mu q}},-_{F_{\mu q}}\rangle$ or $|-_{E_{\mu q}},+_{F_{\mu q}}\rangle$ means finding the state to be $|\zeta^{+}_{mq}\rangle$, and an odd number of odd parity results signifies obtaining $|\zeta^{-}_{EF,mq}\rangle$ as our measurement outcome. Putting together the pieces along with the two cases for $f(\textbf{h}_{mq})$, we obtain the conditional probabilities $p_{\pm}(x_{s})$ that given an incoming photon is in spatial mode $q$, it is in the symmetric or anti-symmetric combinations $|\phi_{AB,q}^{\pm}\rangle$ of the signals collected at the two sites
@@ VI Quantum limit of precision for estimating two-source separation
where $\psi(x)$ is the PSF, and $N$ is the total number of incoming photons during the integration time. Taking the combined PSF (14) for our system of 2 hard apertures, and setting $\beta=\pi r/\sigma$, where $r=2\beta/\delta$ is the ratio between the mid-point separation $2\beta$ (between the two apertures) and the individual aperture size $\delta=2\pi/\sigma$, gives the following QFI
@@
Now, recall that $\sigma=2\pi/\delta$, with $\delta$ the size and $r=2\beta/\delta$. Also, the total number $N$ of photons collected during the integration time will be proportional to the area of the apertures $\delta^{2}$. So all in all, the $4N\pi^{2}/\sigma^{2}$ factor will scale as $\delta^{4}$. The $3r^{2}$ term in the QFI will scale as $\delta^{2}\beta^{2}$, whereas the other term will be proportional to $\delta^{4}$. This tells us that as $\delta$ gets larger (or equivalently as $r$ is small and closer to unity), we will get more information from the spatial mode resolution of the light within each aperture; whereas in the other extreme of $\delta$ being small ($r\gg\beta$), loading the photons contained within the local fundamental (PSF) spatial modes of each telescope’s primary mirror into the quantum memory (i.e., no mode sorter, or $K=1$) will already achieve performance close to the QFI. A fully-general statement to this effect—regarding the relative-information-contributions from the local spatial modes versus the baseline—to more complex quantitative imaging problems such as multi-source localization, moment estimation, object classification, and change detection, etc., are left open for future work.